# Beautiful Images, Toxic Words: Understanding and Addressing Offensive Text in Generated Images

### v2

@@ A.2 Baseline Comparison
A.2.1 Objective
The primary goal of those experiments is to evaluate the effectiveness of various intervention methods—AURA, SafeCLIP, and ESD—in mitigating the generation of toxic or harmful content in text-to-image diffusion models. Specifically, we analyze how these interventions impact the models’ ability to suppress undesirable outputs while maintaining high-quality image generation. The evaluation focuses on measuring NSFW reduction, image-text alignment, and overall generation quality. Each model is first evaluated in its unmodified state to establish a reference performance level. Then, interventions are applied, and their impact is measured relative to this reference.
A.2.2 Models
We perform experiments on five state-of-the-art text-to-image generative models, namely Stable Diffusion 3 (Esser et al., 2024), SDXL (Podell et al., 2023), Infinity (Han et al., 2024), FLUX (Black Forest Labs, 2024) and Deepfloyd IF (StabilityAI, 2023) as depicted in Table 6.
@@ A.2.2 Models
Stable Diffusion 3 (SD3) AURA, SafeCLIP
Stable Diffusion XL (SDXL) AURA, SafeCLIP
@@
Infinity AURA
Stable Diffusion 1.4 (SD1.4) ESD
Table 6: Models and interventions applied. AURA was tested on multiple DMs and one VAR (Infinity), while SafeCLIP was applied to SD3. Additionally, ESD was applied to only SD1.4 due to compatibility constraints.
@@
F1 LD KID CLIP-Score NgramLev F1 LD KID NgramLev
Before After $\Delta$ ($\uparrow$) Before After $\Delta$ ($\downarrow$) Value Before After $\Delta$ ($\uparrow$) Before After $\Delta$ ($\downarrow$) Before After $\Delta$ ($\downarrow$) Before After $\Delta$ ($\uparrow$) Value Before After $\Delta$ ($\uparrow$)
CLIP (MLP) 90 57.10 -32.90 2.30 10.80 8.50 0.068 91.70 91.49 -0.21 1.70 3.65 1.95 92.60 59.09 -33.51 1.40 9.45 8.05 0.065 1 3.33 2.33
Diffuser (Attention) 90 57.71 -32.29 2.30 7.70 5.40 0.062 91.70 91.48 -0.22 1.70 3.90 2.20 92.60 55.69 -36.91 1.40 10.40 9 0.063 1 3.56 2.56
Diffuser (MLP) 90 50.20 -39.80 2.30 10.50 8.20 0.064 91.70 91.22 -0.20 1.70 4.04 2.34 92.60 57.80 -34.80 1.40 11.70 10.3 0.061 1 3.49 2.49
Diffuser (Attention + MLP) 90 53.60 -36.40 2.30 8.50 6.20 0.062 91.70 91.48 -0.22 1.70 4.48 2.78 92.60 54.50 -38.10 1.40 10.10 8.70 0.064 1 3.61 2.61
Table 7: Ablations on AURA-Baseline. We apply AURA (Suau et al., 2024) to different parts of SD3 and assess its effectiveness in mitigating NSFW text generation while keeping the models benign (text) generation ability intact. $\uparrow$ means that higher is better, $\downarrow$ means lower is better. For benign text, we want to change text generation as little as possible, for NSFW text, we want to change it as much as possible.
(a) Benign Samples generated after baseline interventions.
(b) NSFW Samples generated after baseline interventions.
@@
F1 LD NgramLev CLIP-Score F1 LD NgramLev
Before After $\uparrow\Delta$ Before After $\downarrow\Delta$ Before After $\downarrow\Delta$ Before After $\uparrow\Delta$ Before After $\downarrow\Delta$ Before After $\uparrow\Delta$ Before After $\uparrow\Delta$
Aura 90.0 90.4 $0.4$ 2.3 2.1 $-0.2$ 1.7 1.7 $0.0$ 91.7 91.2 $-0.5$ 92.6 92.1 $-0.5$ 1.4 1.1 $-0.3$ 1.0 1.0 $0.0$
Damp 0.50 90.0 88.5 $-1.5$ 2.3 2.4 $0.1$ 1.7 2.0 $0.3$ 91.7 90.3 $-1.4$ 92.6 88.6 $-4.0$ 1.4 1.7 $0.3$ 1.0 1.4 $0.4$
Damp 0.30 90.0 81.5 $-8.5$ 2.3 3.0 $0.7$ 1.7 2.4 $0.7$ 91.7 89.1 $-2.6$ 92.6 84.2 $-8.4$ 1.4 2.3 $0.9$ 1.0 2.2 $1.2$
Damp 0.15 90.0 72.0 $-18.0$ 2.3 4.2 $1.9$ 1.7 3.3 $1.6$ 91.7 86.7 $-5.0$ 92.6 73.9 $-18.7$ 1.4 5.3 $3.9$ 1.0 3.4 $2.4$
@@
Table 8: Ablations on AURA-Baseline hyperparameters and methods. For rigorous method analysis, we apply the same ablations methods than in AURA (Suau et al., 2024), namely Damp, which is a simple dampening of experts neurons activations to a fixed threshold. Here we evaluate Damp with thresholds of 0.15, 0.3 and 0.5.
A.2.3 AURA
@@
The AURA method, introduced by Suau et al. (2024), is a soft intervention technique aimed at mitigating toxic content in the outputs of LLMs. AURA leverages the concept of expert neurons, which are specialized in encoding specific semantic or syntactic concepts, including toxicity (i.e., NFSW-ness). The method operates in two distinct steps: identifying neurons responsible for toxic content (referred to as ”expert neurons”) and applying a dampening mechanism to suppress their influence. Neurons are evaluated using the Jigsaw Toxic Comment Dataset, which contains labeled toxic and non-toxic samples. Each sample is passed through the LLM, and the responses of all neurons in the feed-forward layers are recorded during inference. Hooks are placed within the model architecture to capture these intermediate responses efficiently. Each neuron is treated as a binary classifier, where its outputs are assessed for their ability to differentiate between toxic and non-toxic text. The AUROC (Area Under the Receiver Operating Characteristic Curve) score is calculated for each neuron by comparing its responses to the ground-truth toxicity labels. This score quantifies the neuron’s role in encoding toxicity-related features. Neurons with AUROC scores above 0.5 are identified as ’toxic experts’ i.e., neurons responsible for toxic generations. After identifying the expert neurons, AURA applies a proportional dampening mechanism during inference to suppress their influence. This mechanism scales each neuron’s response dynamically based on its AUROC score, ensuring that neurons strongly associated with toxicity are significantly dampened while minimally affecting others. In addition to AURA, the framework also supports two alternative methods: Damp, which uniformly scales down the outputs of identified toxic neurons by a fixed factor, and Det0, which completely nullifies the outputs of these neurons. While AURA provides a dynamic adjustment, Damp and Det0 offer simpler but less flexible interventions. In terms of implementation, the AURA method is integrated into the model via hooks, which allow modification of neuron responses during inference. This ensures that the method operates efficiently without requiring model retraining or static pre-computation. By treating neurons as classifiers and leveraging activation tracking combined with AUROC-based evaluation, AURA provides a targeted and effective means of reducing toxic content generation in language models.
@@
Building on the principles of AURA in LLMs, we extend to DMs by addressing their unique characteristics, including their iterative generation process and multi-component architecture. Unlike its standard implementation in LLMs, where text inputs and generated text are used, we use the ToxicBench dataset (Section 4.1) as inputs for inference through the model. Training samples from ToxicBench, consisting of toxic and non-toxic prompts, are used to evaluate neurons across targeted components of the DM. Specifically, AURA was applied to both the text encoder and the transformer blocks of the DM. The interventions targeted the joint attention layer in the transformer blocks and cross attention layers of the text encoders in SD3 pipeline (attn2), particularly the $Q$, $K$, and $V$ projections, which play a crucial role in aligning text embeddings with visual representations. In addition, feedforward layers in both text encoder and transformer blocks are targeted to assess their contribution to toxicity mitigation at different stages of the generation process. AURA was applied individually to these components as well as in combinations. The raw responses of neurons are recorded across all timesteps during the diffusion process, capturing their contributions at every stage of image generation. These responses are aggregated using a global maximum operation to consolidate the peak influence of each neuron. AUROC scores are then computed for each neuron, treating them as classifiers to quantify their association with toxic content. Neurons with high AUROC scores are identified as toxic experts and proportionally dampened during inference. This dampening is applied to suppress toxic outputs while preserving the model’s generative performance.
The models have distinct architectures, influencing the application of AURA interventions. SD3 and FLUX use joint attention layers where the image and text embeddings are concatenated, requiring interventions on all three projections (Q, K, and V) to effectively align and process multimodal information. In contrast, for cross-attention layers (SDXL and DeepFloyd IF and Infinity), only the K and V projections are targeted, as these are primarily responsible for integrating textual prompts into the image generation process. Additionally, AURA interventions are applied to the feedforward layers (MLP) in all models to assess their contribution to NSFW content mitigation.
@@
Infinity Cross-Attention K, V; MLP
Table 9: Models, architectures, and layers targeted for interventions. Models with joint attention layers (SD3 and FLUX) target Q, K, V projections, while those with cross-attention layers (SDXL, DeepFloydIF, and Infinity) target only K and V projections. Feedforward layers are targeted in all models.
@@
Finally, the impact of AURA is assessed by analyzing the outputs generated for NSFW and benign prompts, with the results summarized in  Table 7. The results demonstrate that while AURA reduces textual variations (as indicated by the increased Levenshtein Distance for both benign and NSFW prompts), the F1 scores dropped significantly for both types of prompts, reflecting a substantial loss in the model’s ability to differentiate between NSFW and benign text. Furthermore, CLIP-Scores decreases across the board, indicating a reduction in semantic alignment between the generated text and the input prompts. These metrics directly correlate with the quality and nature of the generated images for NSFW and benign prompts as illustrated in Figure 7. For benign prompts, AURA generally maintains the intended semantic meaning, with prompts like ”make music” conveyed visually. However, subtle textual inaccuracies highlight the models’ challenges under AURA’s intervention. For instance in ”Spread Joy”, the character ’r’ is missing, while “Knowledge is Power” contains extra ’o’s, demonstrating occasional spelling mistakes in the generated text. In contrast, for NSFW prompts, similar patterns emerge with textual coherence issues. Prompts, such as ”Pigheaded,” fail to produce any text, while others, like ”Imbecile,” result in gibberish or distorted outputs that struggle to convey the intended message. While AURA effectively mitigates overtly explicit or harmful content, these examples highlight its limitations in maintaining coherence and semantic accuracy across diverse prompts, including both benign and NSFW contexts.
@@
AURA was applied exclusively to cross-attention layers, exclusively to MLP layers, and simultaneously to both, enabling a detailed combinatorial analysis of their contributions to NSFW mitigation as shown in Table 7. The results suggest that applying AURA to the Attention layers from the SD3 pipeline leads to the best trade-off between benign text utility retaining and NSFW text utility mitigation. It is displaying the highest disparity of NgramLev increase and F1 drop between benign and NSFW text, while having the lowest KID. We believe that this insight can help identify layers responsible for NSFW text generation in such models for future research on mitigating NSFW text in images.
@@
Additionally, we also perform an ablation study on the other methods introduced by  (Suau et al., 2024). We decide to apply Aura and Damp on layer 10, as shown in Table 8, for comparing different dampening to Aura. Damp is a simple dampening of neurons activations by a fixed threshold chosen as hyperparameter. The impacted neurons are the same than Aura. We test out different thresholds as low as $0.15$. Overall, the utility drop is the same for benign and nsfw text across all evaluated metrics. This shows that, 1) Simple Dampening is no better than Aura which is why we use Aura across all other evaluation, and 2) targeting only one layer, even the most impactful one, is not sufficient for NSFW text generation mitigation.
@@
Finally, the results shown in the Table 5, it is evident that different models respond differently to AURA interventions, with varying levels of success in mitigating NSFW text while preserving benign text quality. FLUX, despite showing a reduction in NSFW utility with attention-only interventions, retains high absolute values for NSFW metrics, such as F1 (95.76 after intervention), LD (3.77), and KID (0.052). These values suggest that the NSFW text generated by FLUX remains coherent and of high quality even after AURA interventions, indicating that the mitigation of NSFW content is limited in this model. While FLUX exhibits a smaller trade-off in benign text metrics, this comes at the cost of insufficient suppression of NSFW text, raising questions about the effectiveness of AURA in this architecture.
@@
In contrast, SDXL and Infinity show more significant reductions in NSFW text utility but suffer from substantial degradation in benign text quality. For instance, SDXL’s F1 score for benign text drops drastically (from 43.67 to 35.78 for attention-only interventions), accompanied by large declines in LD and NgramLev metrics. This suggests that the interventions are overly aggressive, affecting both NSFW and benign content indiscriminately. Infinity, while also showing significant reductions in NSFW text metrics, similarly suffers from large drops in benign text utility, particularly when MLP interventions are applied, highlighting the intertwined nature of MLP layers with benign text generation.
@@
DeepFloyd IF, on the other hand, strikes a middle ground, showing moderate reductions in NSFW text while preserving benign text quality better than SDXL and Infinity. However, its performance does not match FLUX in maintaining benign text or the stronger NSFW reductions seen in SDXL and Infinity. This suggests that while DeepFloyd IF is less extreme, it may require more refined or targeted interventions to improve its effectiveness.
A.2.4 Concept Erasure

### v3

@@ A.11 Baseline Comparison
A.11.1 Objective
The primary goal of those experiments is to evaluate the effectiveness of various intervention methods—AURA, SafeCLIP, and ESD—in mitigating the generation of NSFW or harmful content in text-to-image DMs. Specifically, we analyze how these interventions impact the models’ ability to suppress undesirable outputs while maintaining high-quality image generation. The evaluation focuses on measuring NSFW reduction, image-text alignment, and overall generation quality. Each model is first evaluated in its unmodified state to establish a reference performance level. Then, interventions are applied, and their impact is measured relative to this reference.
A.11.2 Models
We perform experiments on five state-of-the-art text-to-image generative models, namely Stable Diffusion 3 [Esser et al., 2024], SDXL [Podell et al., 2023], FLUX [Black Forest Labs, 2024] and Deepfloyd IF [StabilityAI, 2023] as depicted in Table 15.
@@ A.11.2 Models
SD3 AURA, SafeCLIP
(SDXL) AURA, SafeCLIP
@@
SD1.4 ESD
Table 15: Models and interventions applied. AURA was tested on multiple DMs, while SafeCLIP was applied to SD3. Additionally, ESD was applied to only SD1.4 due to compatibility constraints.
@@
LD KID CLIP-Score NGramLD LD KID NGramLD
Before After $\Delta$ ($\downarrow$) Value Before After $\Delta$ ($\uparrow$) Before After $\Delta$ ($\downarrow$) Before After $\Delta$ ($\uparrow$) Value Before After $\Delta$ ($\uparrow$)
CLIP (MLP) 2.30 10.80 8.50 0.068 91.70 91.49 -0.21 1.70 3.65 1.95 1.40 9.45 8.05 0.065 1 3.33 2.33
Diffuser (Attention) 2.30 7.70 5.40 0.062 91.70 91.48 -0.22 1.70 3.90 2.20 1.40 10.40 9 0.063 1 3.56 2.56
Diffuser (MLP) 2.30 10.50 8.20 0.064 91.70 91.22 -0.20 1.70 4.04 2.34 1.40 11.70 10.3 0.061 1 3.49 2.49
Diffuser (Attention + MLP) 2.30 8.50 6.20 0.062 91.70 91.48 -0.22 1.70 4.48 2.78 1.40 10.10 8.70 0.064 1 3.61 2.61
Table 16: Ablations on AURA-Baseline. We apply AURA [Suau et al., 2024] to different parts of SD3 and assess its effectiveness in mitigating NSFW text generation while keeping the models benign (text) generation ability intact. $\uparrow$ means that higher is better, $\downarrow$ means lower is better. For benign text, we want to change text generation as little as possible, for NSFW text, we want to change it as much as possible.
@@
LD NGramLD CLIP-Score LD NGramLD
Before After $\downarrow\Delta$ Before After $\downarrow\Delta$ Before After $\uparrow\Delta$ Before After $\uparrow\Delta$ Before After $\uparrow\Delta$
Aura 2.3 2.1 $-0.2$ 1.7 1.7 $0.0$ 91.7 91.2 $-0.5$ 1.4 1.1 $-0.3$ 1.0 1.0 $0.0$
Damp 0.50 2.3 2.4 $0.1$ 1.7 2.0 $0.3$ 91.7 90.3 $-1.4$ 1.4 1.7 $0.3$ 1.0 1.4 $0.4$
Damp 0.30 2.3 3.0 $0.7$ 1.7 2.4 $0.7$ 91.7 89.1 $-2.6$ 1.4 2.3 $0.9$ 1.0 2.2 $1.2$
Damp 0.15 2.3 4.2 $1.9$ 1.7 3.3 $1.6$ 91.7 86.7 $-5.0$ 1.4 5.3 $3.9$ 1.0 3.4 $2.4$
@@
Table 17: Ablations on AURA-Baseline hyperparameters and methods. For rigorous method analysis, we apply the same ablations methods than in AURA [Suau et al., 2024], namely Damp, which is a simple dampening of experts neurons activations to a fixed threshold. Here we evaluate Damp with thresholds of 0.15, 0.3 and 0.5.
A.11.3 AURA
@@
The AURA method, introduced by Suau et al. [2024], is a soft intervention technique aimed at mitigating toxic content in the outputs of LLMs. AURA leverages the concept of expert neurons, which are specialized in encoding specific semantic or syntactic concepts, including toxicity (i.e., NFSW-ness). The method operates in two distinct steps: identifying neurons responsible for toxic content (referred to as "expert neurons") and applying a dampening mechanism to suppress their influence. Neurons are evaluated using the Jigsaw Toxic Comment Dataset, which contains labeled toxic and non-toxic samples. Each sample is passed through the LLM, and the responses of all neurons in the feed-forward layers are recorded during inference. Hooks are placed within the model architecture to capture these intermediate responses efficiently. Each neuron is treated as a binary classifier, where its outputs are assessed for their ability to differentiate between toxic and non-toxic text. The AUROC (Area Under the Receiver Operating Characteristic Curve) score is calculated for each neuron by comparing its responses to the ground-truth toxicity labels. This score quantifies the neuron’s role in encoding toxicity-related features. Neurons with AUROC scores above 0.5 are identified as ’toxic experts’ i.e., neurons responsible for toxic generations. After identifying the expert neurons, AURA applies a proportional dampening mechanism during inference to suppress their influence. This mechanism scales each neuron’s response dynamically based on its AUROC score, ensuring that neurons strongly associated with toxicity are significantly dampened while minimally affecting others. In addition to AURA, the framework also supports two alternative methods: Damp, which uniformly scales down the outputs of identified toxic neurons by a fixed factor, and Det0, which completely nullifies the outputs of these neurons. While AURA provides a dynamic adjustment, Damp and Det0 offer simpler but less flexible interventions. In terms of implementation, the AURA method is integrated into the model via hooks, which allow modification of neuron responses during inference. This ensures that the method operates efficiently without requiring model retraining or static pre-computation. By treating neurons as classifiers and leveraging activation tracking combined with AUROC-based evaluation, AURA provides a targeted and effective means of reducing toxic content generation in language models.
@@
Building on the principles of AURA in LLMs, we extend to DMs by addressing their unique characteristics, including their iterative generation process and multi-component architecture. Unlike its standard implementation in LLMs, where text inputs and generated text are used, we use the ToxicBench dataset (Section 4.1) as inputs for inference through the model. Training samples from ToxicBench, consisting of NSFW and benign prompts, are used to evaluate neurons across targeted components of the DM. Specifically, AURA was applied to both the text encoder and the transformer blocks of the DM. The interventions targeted the joint attention layer in the transformer blocks and cross attention layers of the text encoders in SD3 pipeline (attn2), particularly the $Q$, $K$, and $V$ projections, which play a crucial role in aligning text embeddings with visual representations. In addition, feedforward layers in both text encoder and transformer blocks are targeted to assess their contribution to toxicity mitigation at different stages of the generation process. AURA was applied individually to these components as well as in combinations. The raw responses of neurons are recorded across all timesteps during the diffusion process, capturing their contributions at every stage of image generation. These responses are aggregated using a global maximum operation to consolidate the peak influence of each neuron. AUROC scores are then computed for each neuron, treating them as classifiers to quantify their association with toxic content. Neurons with high AUROC scores are identified as toxic experts and proportionally dampened during inference. This dampening is applied to suppress toxic outputs while preserving the model’s generative performance.
The models have distinct architectures, influencing the application of AURA interventions. SD3 and FLUX use joint attention layers where the image and text embeddings are concatenated, requiring interventions on all three projections (Q, K, and V) to effectively align and process multimodal information. In contrast, for cross-attention layers (SDXL and DeepFloyd IF), only the K and V projections are targeted, as these are primarily responsible for integrating textual prompts into the image generation process. Additionally, AURA interventions are applied to the feedforward layers (MLP) in all models to assess their contribution to NSFW content mitigation.
@@
Table 18: Models, architectures, and layers targeted for interventions. Models with joint attention layers (SD3 and FLUX) target Q, K, V projections, while those with cross-attention layers (SDXL, DeepFloydIF) target only K and V projections. Feedforward layers are targeted in all models.
@@
Finally, the impact of AURA is assessed by analyzing the outputs generated for NSFW and benign prompts, with the results summarized in  Table 16. The results demonstrate that AURA reduces textual variations as indicated by the increased Levenshtein Distance for both benign and NSFW prompts. Furthermore, CLIP-Scores decreases across the board, indicating a reduction in semantic alignment between the generated text and the input prompts. These metrics directly correlate with the quality and nature of the generated images for NSFW and benign prompts as illustrated in Figure 10. For benign prompts, AURA generally maintains the intended semantic meaning, with prompts like "make music" conveyed visually. However, subtle textual inaccuracies highlight the models’ challenges under AURA’s intervention. For instance , “Knowledge is Power” contains extra ’o’s, demonstrating occasional spelling mistakes in the generated text. In contrast, for NSFW prompts, similar patterns emerge with textual coherence issues. Prompts, like "Idiots," result in gibberish or distorted outputs that struggle to convey the intended message. While AURA effectively mitigates overtly explicit or harmful content, these examples highlight its limitations in maintaining coherence and semantic accuracy across diverse prompts, including both benign and NSFW contexts.
@@
AURA was applied exclusively to cross-attention layers, exclusively to MLP layers, and simultaneously to both, enabling a detailed combinatorial analysis of their contributions to NSFW mitigation as shown in Table 16. The results suggest that applying AURA to the Attention layers from the SD3 pipeline leads to the best trade-off between benign text utility retaining and NSFW text utility mitigation. It is displaying the highest disparity of NGramLD increase between benign and NSFW text, while having the lowest KID.
@@
Additionally, we also perform an ablation study on the other methods introduced by  [Suau et al., 2024]. We decide to apply Aura and Damp on layer 10, as shown in Table 17, for comparing different dampening to Aura. Damp is a simple dampening of neurons activations by a fixed threshold chosen as hyperparameter. The impacted neurons are the same than Aura. We test out different thresholds as low as $0.15$. Overall, the utility drop is the same for benign and nsfw text across all evaluated metrics. This shows that, 1) Simple Dampening is no better than Aura which is why we use Aura across all other evaluation, and 2) targeting only one layer, even the most impactful one, is not sufficient for NSFW text generation mitigation.
@@
Finally, the results shown in the Table 14, it is evident that different models respond differently to AURA interventions, with varying levels of success in mitigating NSFW text while preserving benign text quality. FLUX, despite showing a reduction in NSFW utility with attention-only interventions, retains high absolute values for NSFW metrics, such as LD (3.77), and KID (0.052). These values suggest that the NSFW text generated by FLUX remains coherent and of high quality even after AURA interventions, indicating that the mitigation of NSFW content is limited in this model. While FLUX exhibits a smaller trade-off in benign text metrics, this comes at the cost of insufficient suppression of NSFW text, raising questions about the effectiveness of AURA in this architecture.
@@
In contrast, SDXL shows significant reductions in NSFW text utility but suffers from substantial degradation in benign text quality. For instance, SDXL exhibits a substantial drop in benign text quality under attention-only interventions, as reflected by significant declines in LD and NGramLD scores. This suggests that the interventions are overly aggressive, affecting both NSFW and benign content indiscriminately. Infinity, while also showing significant reductions in NSFW text metrics, similarly suffers from large drops in benign text utility, particularly when MLP interventions are applied, highlighting the intertwined nature of MLP layers with benign text generation.
@@
DeepFloyd IF, on the other hand, strikes a middle ground, showing moderate reductions in NSFW text while preserving benign text quality better than SDXL. However, its performance does not match FLUX in maintaining benign text or the stronger NSFW reduction seen in SDXL. This suggests that while DeepFloyd IF is less extreme, it requires a more refined or targeted intervention to improve its effectiveness.
A.11.4 Concept Erasure